OpenStudy (anonymous):

let f(x)=e^-1 and g(x)-1n(x+2). Find all points of intersection for the graphs of f and g.

OpenStudy (anonymous):

need a little clarification here. is it \[f(x)=e^{-1}\] ? that is a constant. and \[g(x)=\ln(x+2)\]? \[g(x)=-\ln(x+2)\]?

OpenStudy (anonymous):

sorry..typo there... it she bef(x)=e^x-1

OpenStudy (anonymous):

shd

OpenStudy (anonymous):

g(x)=ln(x+2)

OpenStudy (anonymous):

\[f(x)=e^{x-1}\] is that it?

OpenStudy (anonymous):

\[f(x)=e ^{x}-1\]

OpenStudy (anonymous):

retype question: Let \[f(x)= e ^{x}-1\] and \[g(x)=1n(x=2)\]. Find all points of intersection for the graphs of f and g.